OpenStudy (anonymous):

Evaluate: ∫x(x³+3)dx

OpenStudy (raden):

u need simplify it be : x(x³+3) = x^4+3x now, integrate

OpenStudy (anonymous):

it then become 4x³+3?

OpenStudy (anonymous):

nvm I'm not suppose to derive right?

OpenStudy (raden):

yes, that is derivative :)

OpenStudy (anonymous):

what is n suppose to be then?

OpenStudy (raden):

oh, u have remembered the basic formula before : int (x^n) dx = 1/(n+1) x^(n+1) look, u can divided 2 cases like this : int (x^4+3x) dx = int x^4 dx + int 3x dx just integrate one by one

OpenStudy (anonymous):

so for int 3x dx, n is 1?

OpenStudy (raden):

yup, or u can move the 3 to infront the sign of integral : int (3x) dx = 3 int x dx

OpenStudy (raden):

so, 3 as the coefficient... just find int x, then times 3

OpenStudy (anonymous):

ok, so for the first int, I got: \[\frac{ 1 }{ 5 }x ^{5}+C\] For second int, I got: \[3\int\limits_{}^{}\frac{ 1 }{ 2 }x ^{2}dx\]

OpenStudy (anonymous):

+C along with the second int

OpenStudy (anonymous):

is that correct?

OpenStudy (raden):

for the 2nd, why still there is integral sign :)

OpenStudy (anonymous):

I thought I have to put the 3 in front of the integral sign, so I left it

OpenStudy (raden):

yes, after u have integrate... miss the sign integral, u have found its result

OpenStudy (anonymous):

so it would be \[3\frac{ 1 }{ 2 }x ^{2}dx\]?

OpenStudy (raden):

just 3/2 x^2 dont use dx again :)

OpenStudy (raden):

3 * 1/2 = 3/2

OpenStudy (anonymous):

Do I have to add C like \[\frac{ 3 }{ 2 }x^{2}+C\]?

OpenStudy (raden):

yes, all solution is 1/5x^5 + 3/2x^2 + c

OpenStudy (raden):

it is indefinite integrals, so always endly + c

OpenStudy (anonymous):

ohh i see, Ty

OpenStudy (raden):

your're welcome